# A method to derive material-specific spin-bath model descriptions of materials displaying prevalent spin physics (for simulation on NISQ devices)

### v2

@@ III.2 Closed configuration of para-benzyne
Figure 3: (a) Molecular structure of the closed configuration of the molecule para-benzyne ($\text{C}_{6}\text{H}_{4}$). (b) Isosurface of one spin-like orbital $\phi_{q=1}$ of para-benzyne. (c) Isosurface of the other spin-like orbital $\phi_{q=2}$ of para-benzyne.
As a second testbed for the parity optimization procedure we have adopted the molecule para-benzyne $\text{C}_{6}\text{H}_{4}$ [19]. The molecule is depicted in Fig. 3(a). In its closed configuration para-benzyne is known to feature two spin-like orbitals that are non-trivial linear combinations of the original basis orbitals [20]. The complete active space self-consistent field (CASSCF) [18, 21, 22] method was used to compute the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ in the CASSCF ground state. The basis set comprises $N_{\phi_{i}}=62$ CASSCF canonical molecular orbitals of which $N_{\phi_{i,a}}=12$ orbitals form the active space of the calculation. The Hamiltonian description of the molecule in the basis reads
@@
where the one-electron integrals $t^{\sigma\sigma^{\prime}}_{ij}$ and two-electron integrals $V^{\sigma\sigma^{\prime}}_{ijkl}$ have been determined as part of an ab-initio calculation. The parity optimization procedure was performed for two sets of initial basis orbitals. The first set contained all $N_{\phi_{i}}=62$ basis orbitals of the CASSCF calculation. The second set was restricted to the $N_{\phi_{i}}=12$ basis orbitals $\phi_{i}$ with local parity $\langle P_{i}\rangle_{0}<9.5\times 10^{-1}$, which coincides with the active space of the CASSCF calculation.
@@ III.3 Single impurity Anderson model
where the magnetic impurity is coupled to the elctrons in the system via its spin degree of freedom $\vec{S}_{d}=(S^{x}_{d},S^{y}_{d},S^{z}_{d})$ only.
@@
We have studied a SIAM Hamiltonian of length $N=9$ system lattice sites and an additional site representing the magnetic impurity with hybridization between system and impurity $V_{0}/t=1/4$. The hybridization $t$ between system lattice sites represents the energy scale of the system. We have fixed the electron number $N_{\mathrm{electrons}}=10$ and $\sum_{i}\sigma^{z}_{i}=0$ to perform a ground state computation using a numerical diagonalization method for a range of impurity interaction strength values $3\leq U/t\leq 40$. We have calculated the local parity $\langle P_{d}\rangle_{0}$ of the impurity site in the ground state. We have used our Schrieffer-Wolff transformation approach to numerically determine the Hamiltonians $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ for each value of the interaction strength $U/t$ and have compared them with the established and analytic results (III.3) and (64). The numerical results match the analytic expressions exactly. We have also calculated the discrepancy $\Delta E_{0}=|E_{0,(\mathcal{P})\tilde{H}(\mathcal{P})}-E_{0,H}|$ between the ground state energies $E_{0}$ of the SIAM Hamiltonian $H$ (III.3) and the energy of respective ground states of the transformed Hamiltonians $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$. The ground state local parity $\langle P_{d}\rangle_{0}$ of the magnetic impurity and the rescaled ground energy discrepancies $U\Delta E_{0}/t$ as a function of the interaction strength $U/t$ are shown in Fig. 5.
@@
Figure 5: Rescaled ground state energy difference $U\times\Delta E_{0}/t=U\times|E_{0,(\mathcal{P})\tilde{H}(\mathcal{P})}-E_{0,H}|/t$ between the ground state of the SIAM Hamiltonian and the transformed Hamiltonians $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ as a function of the interaction strength $U/t$. The dark grey line represents the scaled energy difference for the transformed Hamiltonian $\tilde{H}$. The dashed orange displays $U\times\Delta E_{0}/t$ for the effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ where the impurity site is restricted to single occupancy $n_{d}\equiv 1$. The solid blue line displays the local parity $\langle P_{d}\rangle_{0}$ of the impurity site in the true ground state for each value of the interaction strength $U/t$.
@@
For values of the interaction strength $U/t\leq 4$, we observe a pronounced energy difference between the SIAM ground state and the respective ground states of the transformed and effective Hamiltonians. In this weak interaction regime, the chemical potential of the impurity lies within the bandwith $D=4t$ of the tight-binding chain representing the remainder of the system. There is then no energy gap between the subspace of the Hilbert space where the impurity is singly occupied and the rest of the Hilbert space. In the absence of an energy gap, limiting the series expansion (18) to first order in $S$ is not a justified approximation. For values of the interaction strength $U/t>4$ we observe a decrease of the energy discrepancy $\Delta E_{0}$ of the ground state energies of the SIAM Hamiltonian and the transformed and effective Hamiltonians respectively as a power-law $(U/t)^{-\alpha}$ of the interaction strength with an exponent $\alpha>2$. The energy discrepancy quickly becomes small with respect to the energy scale $t$ and we find $\Delta E_{0}/t\rightarrow 0$ for $U/t\rightarrow\infty$. For the ground state local parity on the impurity site we observe $7\times 10^{-2}>(\langle P_{d}\rangle_{0}+1)>3\times 10^{-2}$ in the regime $U/t\leq 4$. For larger values of the interaction strength $U/t>4$ we find that the local parity of the impurity site approximates a polynomial function $(\langle P_{d}\rangle_{0}+1)(U/t)\propto(U/t)^{-\beta}$ with the exponent $\beta\approx\alpha-1$. The local parity of the impurity site appears to be a good indicator of quality of the approximation (19) if $\tilde{H}$ is supposed to become block-diagonal in the subspaces of the Hilbert space with different particle number $n_{d}$ on the impurity site. This supports our claim that a local parity $(\langle P_{d}\rangle_{0}+1)\rightarrow 0$ signals that the corresponding site can be represented as a spin degree of freedom and that $\mathcal{P}\tilde{H}\mathcal{P}$ is then a good effective Hamiltonian to describe the low-energy spin dynamics of the system. The energy discrepancy $\Delta E_{0}$ is identical for both $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$. In the case of the SIAM Hamiltonian the Schrieffer-Wolff transformation yields a Hamiltonian $\tilde{H}$ that does no longer couple the distinct subspaces characterized by $n_{d}=1$ and $n_{d}\neq 1$. We find that our Schrieffer-Wolff transformation method recovers the analytically known results and that the ground state of effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ approaches the true ground state as $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-1}$.
@@ III.4 Disordered Fermi-Hubbard model
Another significant model Hamiltonian, for which the analytic expression for the Schrieffer-Wolff transformed Hamiltonian is known, is the Fermi-Hubbard model [27, 28]. It is translation invariant and describes the electrons in orbitals with small nearest-neighbor hybridizations $t$ where electrons occupying the same orbital exert a density-density interaction $U$ on each other. In the limit $U/t\rightarrow\infty$ and $N_{\mathrm{electrons}}/N_{\mathrm{orbitals}}=1$ the transformed Hamiltonian $\tilde{H}$ reduces to the well-known Heisenberg model of interacting spins [16, 12]. Here, we present the Schrieffer-Wolff transformation for the disordered Fermi-Hubbard (dFH) chain. Its Hamiltonian is given by
@@
For values of the interaction strength $U/t\leq 10$, we observe a significant energy difference $\Delta E_{0}$ for both $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ which is decreasing with $U/t$ at an increasing rate faster than $(U/t)^{-1}$. At values $U/t>10$ the energy difference becomes small with respect to the energy scale $t$ and is approaching $\Delta E_{0}\rightarrow 0$ for $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-2}$. Similarly we find average local parity $(\overline{\langle P_{i}\rangle_{0}}+1)>10^{-1}$ for $U/t\leq 10$. This highlights that the chain sites cannot be considered good realizations of spin degrees of freedom in this regime. The transformed Hamiltonian $\tilde{H}$ can consequently not be considered a good approximation to the Hamiltonian $H$ (65). This is supported by the respective values of $\Delta E_{0}/t\geq 1$. As $(\overline{\langle P_{i}\rangle_{0}}+1)(U/t)\ll 10^{-1}$ for $U/t>10$ we also observe that the transformed Hamiltonian becomes an ever better approximation of the original dFH Hamiltonian and $\Delta E_{0}/t\ll 1$. For the disordered Fermi-Hubbard Hamiltonian we again find that our Schrieffer-Wolff transformation method recovers the analytically known results. The ground state of effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ approaches the true ground state as $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-1}$ and the crossover at which $\mathcal{P}\tilde{H}\mathcal{P}$ becomes a good description of the low-energy dynamics of the model is characterized by $(\overline{\langle P_{i}\rangle_{0}}+1)<10^{-1}$
@@ III.5 Molecular Chromium bromide
An interesting testbed for our spin mapping approach, is the class of 2D magnetic materials, which have recently been discovered experimentally[29, 30]–seemingly in contradiction to the Mermin-Wagner theorem. A common approach to study 2D magnetic materials from first principles is to perform a so-called “DFT+U” calculation, followed by a Wannierization and the extraction of a classical Heisenberg-like spin model from the resulting tight-binding description of the material (cf. [31] for a recent study on Chromium tri-halides). The spin mapping proposed in this work represents an alternative for extracting the effective spin model. The key difference to the aforementioned workflow to extract effective spin models using DFT is: 1) A correlated quantum-chemistry (wave-function) approach is used. 2) The Hamiltonian resulting from the Schrieffer-Wolff transformation is still a quantum mechanical Hamiltonian, i.e., the classical limit is not implied by the spin mapping approach. Let us point out that we neither claim that a full quantum mechanical treatment of the effective spin model is required for getting an accurate spin model for (ferromagnetic) transition-metal tri-halides, nor that a correlated method is required for the electronic-structure calculation (however, standard DFT, without “+U”, is not sufficient to capture the localized Chromium $d$-orbitals). We would also like to note that the main purpose of the multi reference calculation is the determination of the spin orbitals, i.e. the single particle basis. We can use hints from multi reference methods to restrict the application of the Schrieffer-Wolff transformation to a subset of all orbitals, but this step is not essential.
@@
We initially perform a series of electronic structure calculations using the def2-QZVP basis set to generate the information necessary for a search of the active space of orbital basis of the system. We have used the Active Space Finder (ASF), [32] an open source package to assist the user in the selection of an active space. As expected, five Chromium $d$-orbitals and $p$-orbitals from the outer Shell of Bromide are suggested by the ASF and the active space contains $N_{\phi_{i}}=9$ basis orbitals. We have performed a CASSCF calculation using the previously identified orbitals as the active space to compute the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ for the CASSCF ground state. From this CASSCF calculation we have also obtained the one-electron $t^{\sigma\sigma^{\prime}}_{ij}$ and two-electron integrals $V^{\sigma\sigma^{\prime}}_{ijkl}$ in the active space. We can write the Hamiltonian description of the active space of chromium bromide as
@@
where $c^{\dagger}_{i\sigma}$ creates an electron with spin $\sigma$ in the orbital $\phi_{i}$ of the active space. Using the reduced density matrices we have performed a parity optimization of the basis orbitals of the active space. Here, we first ran a parity minimization procedure of the entire basis and identified $N_{\phi_{q}}=3$ spin-like orbitals with respective parities $\langle P_{q}\rangle_{0}\leq-9.94\times 10^{-1}$. In a subsequent optimization run have maximized the local parities of the remaining basis orbitals. The local parities of the initial basis orbitals $\phi_{i}$ and the optimized basis orbitals $\phi_{q}$ are shown together with the average electron density $\langle n_{q}\rangle_{0}$ in Fig. 7.
@@
where $\dots$ denotes minor contributions from other basis orbitals $\phi_{i}$.
[…]
@@
[…]
resulting from the Schrieffer-Wolff transformation procedure comprises $N_{h}=1.13\times 10^{6}$ terms of which roughly $N_{|h|\geq c}=22968$ terms feature a coupling constant $|h|\geq c$. The projection of the transformed Hamiltonian to the subspace of the Hilbert space, in which the spin-like orbitals $\phi_{q\in\mathcal{S}}$ are replaced with spin degrees of freedom, is given by
@@
We have calculated the low energy spectrum of $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ again using a numerical diagonalization method and with the same constraints for the quantum numbers. The low-energy spectra of the five different numerical diagonalization calculations are displayed in Fig. 8 (a).
@@
Figure 7: Local parities of the active space basis orbitals of molecular chromium bromide $\mathrm{Cr}\mathrm{Br}^{3+}_{6}$ in the CASSCF ground state. The dark grey line displays the local parities of the CASSCF canonical molecular orbitals, which form the active space of the system. The dashed orange line indicates the local parities of the optimized basis orbitals $\phi_{q}$ and the blue circles show the average local electrons density of basis orbitals $\phi_{q}$. The green triangles indicate the average local electron density of the original basis orbitals.
@@
Figure 8: Low energy spectra of the active space Hamiltonians of molecular chromium bromide. (a) The dark grey line shows the eigenenergies of the $n=20$ lowest eigenstates of the original active space Hamiltonian $H$. We display the corresponding eigenenergies of the effective Hamiltonian $\mathcal{P}H\mathcal{P}$ (solid blue), projected onto the subspace of the Hilbert space where the spin-like orbitals are restricted to single occupancy, the transformed Hamiltonian $\tilde{H}$ (dashed orange), and the effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ (dashed green) acting on the subspace $\mathcal{P}$ in which the fermionic degrees of freedom in the spin-like orbitals have been substituted with spin degree of freedom. The low energy spectrum of the effective Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$, acting only on the subspace characterized by $n_{i}\equiv 1$, for $i\in\{4,6,7\}$, is shown in purple. The average local parity of the spin-like orbitals $\overline{\langle P_{q\in\mathcal{S}}\rangle}_{n}$ for the lowest $n=20$ eigenstates $|n\rangle$ of the original Hamiltonian $H$ is displayed as grey circles. (b) Low energy spectra reduced to the eigenstates $|n_{\mathcal{S}}\rangle$ of $H$ and $\tilde{H}$, which satisfy $\overline{\langle P_{q\in\mathcal{S}}\rangle}_{n_{\mathcal{S}}}\simeq-1$.

### v3

@@ III.2 Closed configuration of para-benzyne
Figure 3: (a) Molecular structure of the closed configuration of the molecule para-benzyne ($\text{C}_{6}\text{H}_{4}$). (b) Isosurface of one spin-like orbital $\phi_{q=1}$ of para-benzyne. (c) Isosurface of the other spin-like orbital $\phi_{q=2}$ of para-benzyne. The colors red and blue indicate the sign of the orbital wavefunction.
As a second testbed for the parity optimization procedure we have considered the molecule para-benzyne $\text{C}_{6}\text{H}_{4}$ [35]. The molecule is depicted in Fig. 3(a). In its singlet configuration para-benzyne is known to feature two spin-like orbitals that are non-trivial linear combinations of the original basis orbitals [36]. The complete active space self-consistent field (CASSCF) [37, 38, 33, 39] method was used to compute the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ in the ground state. CASSCF wave function is constructed as a linear combination of all Slater determinants obtained by distribution of N electrons into M orbitals, which constitute the (N, M) active space, that is $N$ particles on $M$ spinful orbitals are treated within their full Hilbert space, while the remaining orbitals are treated using a self consistent field approximation . Energy is obtained variationally, being minimized wrt. both active orbitals and Slater determinant coefficients. The orbital basis set comprises $N_{\phi_{i}}=62$ CASSCF canonical molecular orbitals of which $N_{\phi_{i,a}}=12$ orbitals are selected as active containing 12 electrons, giving rise to (12, 12) active space. The Hamiltonian description of the molecule in the molecular orbital basis reads
@@
where the one-electron integrals $t^{\sigma\sigma^{\prime}}_{ij}$ and two-electron integrals $V^{\sigma\sigma^{\prime}}_{ijkl}$ have been determined as part of an ab-initio calculation. The parity optimization procedure was performed for two sets of initial basis orbitals. The first set contained all $N_{\phi_{i}}=62$ basis orbitals of the CASSCF calculation. The second set was restricted to the active $N_{\phi_{i}}=12$ orbitals $\phi_{i}$ with local parity $\langle P_{i}\rangle_{0}<9.5\times 10^{-1}$, which coincides with the active space of the CASSCF calculation.
@@ III.3 Single impurity Anderson model
where the magnetic impurity is coupled to the electrons in the lead via its spin degree of freedom $\vec{S}_{d}=(S^{x}_{d},S^{y}_{d},S^{z}_{d})$ only.
@@
We have studied a SIAM Hamiltonian of with 9 lattice sites for the lead and an additional site representing the magnetic impurity and a hybridization between lead and impurity $V_{0}/t=1/4$. The hybridization $t$ between lead lattice sites represents the energy scale of the system. We have fixed the electron number $N_{\mathrm{electrons}}=10$ and $\sum_{i}\sigma^{z}_{i}=0$ to perform a ground state computation using a numerical diagonalization method for a range of impurity interaction strength values $3\leq U/t\leq 40$. We have calculated the local parity $\langle P_{d}\rangle_{0}$ of the impurity site in the ground state. We have used our Schrieffer-Wolff transformation approach to numerically determine the Hamiltonians $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ for each value of the interaction strength $U/t$ and have compared them with the established and analytic results (III.3) and (64). The numerical results match the analytic expressions exactly. We have also calculated the discrepancy $\Delta E_{0}=|E_{0}^{(\mathcal{P})\tilde{H}(\mathcal{P})}-E_{0}^{H}|$ between the ground state energies $E_{0}$ of the SIAM Hamiltonian $H$ (III.3) and the energy of respective ground states of the transformed Hamiltonians $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$. The ground state local parity $\langle P_{d}\rangle_{0}$ of the magnetic impurity and the rescaled ground energy discrepancies $U\Delta E_{0}/t$ as a function of the interaction strength $U/t$ are shown in Fig. 5.
@@
Figure 5: Rescaled ground state energy difference $U\times\Delta E_{0}/t=U\times|E_{0}^{(\mathcal{P})\tilde{H}(\mathcal{P})}-E_{0}^{H}|/t$ between the ground state of the SIAM Hamiltonian and the transformed Hamiltonians $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ as a function of the interaction strength $U/t$. The dark grey line represents the scaled energy difference for the transformed Hamiltonian $\tilde{H}$. The dashed orange displays $U\times\Delta E_{0}/t$ for the effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ where the impurity site is restricted to single occupancy $n_{d}\equiv 1$. The solid blue line displays the local parity $\langle P_{d}\rangle_{0}$ of the impurity site in the true ground state for each value of the interaction strength $U/t$.
@@
For values of the interaction strength $U/t\leq 4$, we observe several pronounced energy differences between the SIAM ground states and the ground states of the transformed and effective Hamiltonians. In this weak interaction regime, the chemical potential of the impurity lies within the bandwidth $D=4t$ of the tight-binding chain representing the remainder of the system. There is then no energy gap between the subspace of the Hilbert space where the impurity is singly occupied and the rest of the Hilbert space. In the absence of an energy gap, limiting the series expansion (18) to first order in $S$ is not a justified approximation. For values of the interaction strength $U/t>4$ we observe a decrease of the energy discrepancy $\Delta E_{0}$ of the ground state energies of the SIAM Hamiltonian and the transformed and effective Hamiltonians respectively as a power-law $(U/t)^{-\alpha}$ of the interaction strength with an exponent $\alpha>2$. The energy discrepancy quickly becomes small with respect to the energy scale $t$ and we find $\Delta E_{0}/t\rightarrow 0$ for $U/t\rightarrow\infty$. For the ground state local parity on the impurity site we observe $7\times 10^{-2}>(\langle P_{d}\rangle_{0}+1)>3\times 10^{-2}$ in the regime $U/t\leq 4$. For larger values of the interaction strength $U/t>4$ we find that the local parity of the impurity site approximates a polynomial function $(\langle P_{d}\rangle_{0}+1)(U/t)\propto(U/t)^{-\beta}$ with the exponent $\beta\approx\alpha-1$. The local parity of the impurity site appears to be a good indicator of quality of the approximation (19) if $\tilde{H}$ is supposed to become block-diagonal in the subspaces of the Hilbert space with different particle number $n_{d}$ on the impurity site. This supports our claim that a local parity $(\langle P_{d}\rangle_{0}+1)\rightarrow 0$ signals that the corresponding site can be represented as a spin degree of freedom and that $\mathcal{P}\tilde{H}\mathcal{P}$ is then a good effective Hamiltonian to describe the low-energy spin dynamics of the system. The energy discrepancy $\Delta E_{0}$ is identical for both $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$. In the case of the SIAM Hamiltonian the Schrieffer-Wolff transformation yields a Hamiltonian $\tilde{H}$ that does no longer couple the distinct subspaces characterized by $n_{d}=1$ and $n_{d}\neq 1$. We find that our Schrieffer-Wolff transformation method recovers the analytically known results and that the ground state of effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ approaches the true ground state as $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-1}$.
@@ III.4 Disordered Fermi-Hubbard model
Another significant model Hamiltonian, for which the analytic expression for the Schrieffer-Wolff transformed Hamiltonian is known, is the Fermi-Hubbard model [44, 45]. It is translation invariant and describes the electrons in orbitals with small nearest-neighbor hybridizations $t$ where electrons occupying the same orbital exert a density-density interaction $U$ on each other. In the limit $U/t\rightarrow\infty$ and $N_{\mathrm{electrons}}/N_{\mathrm{orbitals}}=1$ (half filling) the transformed Hamiltonian $\tilde{H}$ reduces to the well-known Heisenberg model of interacting spins [16, 11]. Here, we present the Schrieffer-Wolff transformation for the disordered Fermi-Hubbard (dFH) chain. Its Hamiltonian is given by
@@
For values of the interaction strength $U/t\leq 10$, we observe a significant energy difference $\Delta E_{0}$ for both $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ which is decreasing with $U/t$ at an increasing rate faster than $(U/t)^{-1}$. At values $U/t>10$ the energy difference becomes small with respect to the energy scale $t$ and is approaching $\Delta E_{0}\rightarrow 0$ for $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-2}$. Similarly we find average local parity $(\overline{\langle P_{i}\rangle_{0}}+1)>10^{-1}$ for $U/t\leq 10$. This highlights that the chain sites cannot be considered good realizations of spin degrees of freedom in this regime. The transformed Hamiltonian $\tilde{H}$ can consequently not be considered a good approximation to the Hamiltonian $H$ (65). This is supported by the respective values of $\Delta E_{0}/t\geq 1$. As $(\overline{\langle P_{i}\rangle_{0}}+1)(U/t)\ll 10^{-1}$ for $U/t>10$ we also observe that the transformed Hamiltonian becomes an ever better approximation of the original dFH Hamiltonian and $\Delta E_{0}/t\ll 1$. For the disordered Fermi-Hubbard Hamiltonian we again find that our Schrieffer-Wolff transformation method recovers the analytically known results. The ground state of effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ approaches the true ground state as $U/t\rightarrow\infty$ at a rate faster than $(U/t)^{-1}$ and the crossover at which $\mathcal{P}\tilde{H}\mathcal{P}$ becomes a good description of the low-energy dynamics of the model is characterized by $(\overline{\langle P_{i}\rangle_{0}}+1)<10^{-1}$. We note that the Schrieffer Wolff transformation renormalizes the local disorder potentials.
@@ III.5 Molecular Chromium bromide
An interesting testbed for our spin mapping approach, is the class of 2D magnetic materials, which have recently been discovered experimentally[46, 47]–seemingly in contradiction to the Mermin-Wagner theorem. A common approach to study 2D magnetic materials from first principles is to perform a so-called “DFT+U” calculation, followed by a Wannierization and the extraction of a classical Heisenberg-like spin model from the resulting tight-binding description of the material (cf. [48] for a recent study on Chromium tri-halides). The spin mapping proposed in this work represents an alternative for extracting the effective spin model. The key difference to the aforementioned workflow to extract effective spin models using DFT is: 1) A correlated quantum-chemistry (wave-function) approach is used to determine a basis set featuring spin-like orbitals. 2) The Schrieffer–Wolff transformation starts from full description not relying on the approximate multi-reference method used to determine the spin-like orbitals. 3) The Hamiltonian resulting from the Schrieffer-Wolff transformation is still a quantum mechanical Hamiltonian, i.e., the classical limit is not implied by the spin mapping approach.
Let us point out that we neither claim that a full quantum mechanical treatment of the effective spin model is required for getting an accurate spin model for (ferromagnetic) transition-metal tri-halides, nor that a correlated method is required for the electronic-structure calculation (however, standard DFT, without “+U”, is not sufficient to capture the localized Chromium $d$-orbitals). We would also like to note that the main purpose of the multi reference calculation is the determination of the spin orbitals, i.e. the single particle basis. We can use hints from multi reference methods to restrict the application of the Schrieffer-Wolff transformation to a subset of all orbitals, but this step is not essential.
@@
We initially perform a series of electronic structure calculations using the def2-QZVP basis set to generate the information necessary for a search of the active space of orbital basis of the system. We have used the Active Space Finder (ASF), [49] an open-source to assist the user in the selection of an active space. As expected, five Chromium $d$-orbitals and $p$-orbitals from the outer Shell of Bromide are suggested by the ASF and the active space contains $N_{\phi_{i}}=9$ basis orbitals. We have performed a CASSCF calculation using the previously identified orbitals as the active space to compute the reduced density matrices $\rho^{(1)}$ and $\rho^{(2)}$ for the CASSCF ground state. From this CASSCF calculation we have also obtained the one-electron $t^{\sigma\sigma^{\prime}}_{ij}$ and two-electron integrals $V^{\sigma\sigma^{\prime}}_{ijkl}$ in the active space. We can write the Hamiltonian description of the active space of chromium bromide as
@@
where $c^{\dagger}_{i\sigma}$ creates an electron with spin $\sigma$ in the orbital $\phi_{i}$ of the active space. Using the reduced density matrices we have performed a parity optimization of the basis orbitals of the active space. Here, we first ran a parity minimization procedure of the entire basis and identified $N_{\phi_{q}}=3$ spin-like orbital with respective parities $\langle P_{q}\rangle_{0}\leq-9.94\times 10^{-1}$. In a subsequent optimization run have maximized the local parities of the remaining basis orbitals. The local parities of the initial basis orbitals $\phi_{i}$ and the optimized basis orbitals $\phi_{q}$ are shown together with the average electron density $\langle n_{q}\rangle_{0}$ in Fig. 7.
@@
where $\dots$ denotes minor contributions from the other basis orbitals $\phi_{i}$.
We have performed a first numerical diagonalization computation of the $n=20$ lowest energy eigenstates of the effective Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$, which acts exclusively on the subspace of the Hilbert space characterized by $n_{i}\equiv 1$, for $i\in\{4,6,7\}$. In the calculation we have fixed $N_{\mathrm{electrons}}=9$ and $\sum_{i}\sigma^{z}_{i}=3$. Subsequently, the Hamiltonian $H$ was transformed to the parity-optimized orbital basis and a second numerical diagonalization with the same constraints for electron number and total spin $\sigma^{z}$ was performed. For each of the $n=20$ lowest energy eigenstates of the Hamiltonian $H$ we have computed the local parity $\langle P_{q\in\mathcal{S}}\rangle_{n}$ of the orbitals initially identified as spin-like $\phi_{q\in\mathcal{S}}$, where $\mathcal{S}=\{0,1,2\}$. We have projected the Hamiltonian $H$ in the parity-optimized basis to the subspace of the Hilbert space where the number of electrons in the spin-like orbitals $\phi_{q\in\mathcal{S}}$ is fixed $n_{q\in\mathcal{S}}\equiv 1$. Using the same numerical diagonalization method we have computed the $n=20$ lowest energy eigenstates of this projected Hamiltonian $\mathcal{P}H\mathcal{P}$. For the Schrieffer-Wolff transformation we have introduced a cutoff $c=10^{-5}\,\mathrm{Hartree}$ for the constituent terms of the Hamiltonian. Hamiltonian terms of coupling constant $|h|$ smaller than the cutoff have been neglected in the calculation of the transformed Hamiltonian $\tilde{H}$ to limit the memory consumption of the computation. The singular value decomposition in our Schrieffer-Wolff transformation approach was performed using a Krylov subspace method with a maximum dimension of the subspace $\mathrm{dim}(\mathcal{K})=4000$. The size of the vector spaces $\mathcal{V}$ depends on the choice of the cutoff. For chromium bromide we determine $\mathrm{dim}(\mathcal{V}_{0})(c)=8760$ for the initial vector space and $\mathrm{dim}(\mathcal{V}_{1})(c)=3.6\times 10^{6}$ for the target vector space. The prevalent interaction terms in the transformed Hamiltonian read
@@
$+\dots\,$,
and amount to $N_{|h|>c}=41239$ terms with a coupling constant $|h|>c$. The projection of the transformed Hamiltonian to the subspace of the Hilbert space, in which the spin-like orbitals $\phi_{q\in\mathcal{S}}$ are replaced with spin degrees of freedom, is given by
@@
We have calculated the low energy spectrum of $\tilde{H}$ and $\mathcal{P}\tilde{H}\mathcal{P}$ using a numerical diagonalization method and with the same set of constraints for the quantum numbers. The low-energy spectra from numerical diagonalization calculations of the five different Hamiltonian representations are displayed in Fig. 8 (a).
@@
Figure 7: Local parities of the active space basis orbitals of molecular chromium bromide $\mathrm{Cr}\mathrm{Br}^{3-}_{6}$ in the CASSCF ground state. The dark grey line displays the CASSCF ground state local parities of the CASSCF canonical molecular orbitals, which form the active space of the system. The dashed orange line indicates the ground state local parities of the optimized basis orbitals $\phi_{q}$ and the orange circles show the expectation value for the local electrons density of basis orbitals $\phi_{q}$ in the ground state. The dark gray triangles indicate the expectation values for local electron density in the original CASSCF basis orbitals. In the optimized basis there exist three basis orbitals that feature local parities $\langle P_{q}\rangle_{0}<-0.994$, indicating good realizations of three spin degrees of freedom.
@@
Figure 8: Low energy spectra of the active space Hamiltonians of molecular chromium bromide $\mathrm{Cr}\mathrm{Br}^{3-}_{6}$. The dark grey line shows the eigenenergies of the $n=20$ lowest eigenstates of the original active space Hamiltonian $H$. We display the corresponding eigenenergies of the effective Hamiltonian $\mathcal{P}H\mathcal{P}$ (solid blue), projected onto the subspace of the Hilbert space where the spin-like orbitals are restricted to single occupancy, the transformed Hamiltonian $\tilde{H}$ (dashed orange), and the effective Hamiltonian $\mathcal{P}\tilde{H}\mathcal{P}$ (dashed green) acting on the subspace $\mathcal{P}$ in which the fermionic degrees of freedom in the spin-like orbitals have been substituted with spin degree of freedom. The low energy spectrum of the effective Hamiltonian $\mathcal{P}_{0}H\mathcal{P}_{0}$, acting only on the subspace characterized by $n_{i}\equiv 1$ for the CASSCF canonical orbitals $i\in\{4,6,7\}$, is shown in purple. The total local parity of the spin-like orbitals $\sum_{q\in\mathcal{S}}\langle P_{q}\rangle_{n}$ measured for the lowest $n=20$ eigenstates $|n\rangle$ of the original fermionic Hamiltonian $H$ is displayed as grey circles.
Figure 9: (a) Low energy spectra of the respective Hamiltonians reduced to the set of eigenstates $|n_{\mathcal{S}}\rangle$ of $H$ and $\tilde{H}$, for which the expectation value of the total local parity for the spin-like orbitals $\phi_{q}$ satisfies $\sum_{q\in\mathcal{S}}\langle P_{q}\rangle_{n_{\mathcal{S}}}\simeq-3$. (b) Energy deviation $|\delta E_{n_{\mathcal{S}}}|$ between the eigenergies $E_{n_{\mathcal{S}}}$ of the full fermionic Hamiltonian and the respective transformed and projected, effective Hamiltonians. We observe that the spectrum of spinful excitations is better captured by the transformed Hamiltonians $\mathcal{P}\tilde{H}\mathcal{P}$ as opposed to the original Hamiltonian projected onto the subspace of the Hilbert space, where the spin-like orbitals become spin degrees of freedom, $\mathcal{P}H\mathcal{P}$. The energy discrepancy between $H$ and $\mathcal{P}\tilde{H}\mathcal{P}$ remains $|\delta E_{n_{\mathcal{S}}}|\approx 1\times 10^{-3}\,\mathrm{Hartree}$ for the studied excited states. The corresponding discrepancy for the effective Hamiltonian $\mathcal{P}H\mathcal{P}$ is noticeably larger and amounts to $|\delta E_{n_{\mathcal{S}}}|\leq 8\times 10^{-3}\,\mathrm{Hartree}$.